OpenStudy (anonymous):

Solve the triangle completely (2 answers): B= 38.9 degrees, a= 42.y, b= 30

OpenStudy (anonymous):

Sorry, typo. a= 42.7

OpenStudy (anonymous):

use the equations a/sin A = b/sin B= c/sin C or the equation sin A/a = sin B/b= sin C/c

OpenStudy (anonymous):